# Direct Quantized Training of Language Models with Stochastic Rounding

### v2

@@ 4.5 Evaluation Results
In this section, we conduct evaluation on the WikiText-2 test set Merity et al. (2016) and several other zero-shot tasks. We focus on 1B parameter models to best demonstrate the capability of our approach. The results are summarized in Table 1. For consistency, we set the sequence length to 512 across all tasks, matching the configuration used during training. In the table, ‘ternary Inf.’ refers to DQT models trained with 8-bit weights but evaluated using ternary weights. Details on how ternary inference is implemented for DQT models are provided in Appendix A.2. As shown in Table 1, FP32 models achieve the highest overall performance across tasks, followed closely by our DQT 8-bit variants. Notably, except for the WinoGrande task using 1B models trained on the Wikipedia dataset, our DQT 8 bit models outperform BitNet b1.58 across all other benchmarks. These results highlight that DQT 8 bit models more closely approximate the performance of FP32 models compared to BitNet. Moreover, when ternary inference is applied, the performance slightly decreases compared to 8-bit inference but remains on par with BitNet, demonstrating the robustness of our approach with ternary inference and its flexibility in deployment.
@@ 5 Analysis
While stochastic rounding offers advantages in keeping low-precision weights, it may bring some problems. For example, small weight updates may be applied with low probability (e.g., a value like -1+0.02 has only a 2% chance of being rounded to 0). If such rare updates do occur, we are particularly interested in understanding their impact on the overall training dynamics. In this section, we provide a detailed analysis of the effects and implications of using stochastic rounding.
@@ 5.1 The Role of Stochastic Rounding in DQT
Figure 5: Comparison between DQT 1.58 bits and a variant using absmax quantization for weight updates. Both methods use the same learning rate.
Firstly, we provide empirical evidence for the critical role of stochastic rounding in preserving gradient information during training with low-precision weights. Particularly, we compare 130M ternary DQT with a variant that simply uses absmax quantization on the updated weight matrices and maintains the weights in ternary format without stochastic rounding. As shown in Figure 5, the latter fails to converge, despite operating under the same bit budget. This outcome is expected, as absmax quantization can easily ignore small updates on the weights. In contrast, stochastic rounding not only performs quantization but also facilitates the accumulation of fine-grained updates, enabling effective training even in extremely low-bit regimes.
@@ 5.2 Quantifying Weight Update Frequency in DQT
Figure 6: Percentage of updated weights after each training step.
@@
We quantify the weight update frequency in Figure 6 for three variants: DQT 1.58 bit, BitNet b1.58 and DQT 8 bit under the same learning rate and batch size. Note that we show the average percentages of all weight matrices in the model. For BitNet, the peak weight update rate is approximately 0.05% (observed at step 2000, the end of the warm-up phase), meaning only 0.05% of quantized ternary weights change after a single step. Ternary DQT exhibits a similar update rate of around 0.04%, indicating minimal difference between the two in terms of update frequency. In contrast, the 8-bit DQT variant, with weight values ranging from -128 to 127, shows a significantly higher update frequency, reaching up to 8%. A more detailed discussion on how we acquire these percentages is provided in Appendix A.4.
@@ 5.3 Impact of Small Weight Updates in DQT Training
Finally, to investigate the impact of small weight updates during training, we rank the absolute value of weight updates at each step and selectively intervene in the lowest 20%. For these bottom 20% updates, we apply one of two interventions: either suppress the update by retaining the original ternary value, or enforce a change by rounding it to a different quantized value, even if the update is small. For example, if an update of +0.02 from -1+0.02 falls within the smallest 20%, we either keep it at -1 (suppress) or round it to 0 (enforce), depending on the experimental condition. Figure 7 shows the training loss for these variants, where force to remain refers to suppressing the bottom 20% of updates, and force to update corresponds to enforcing them to change. From Figure 7, we observe that the original implementation of DQT (1.58-bit) achieves the best performance. Ignoring the smallest 20% of updates has minimal impact, while enforcing these small updates to take effect slightly accelerates convergence. However, the final losses at the end of training are similar across all variants. This may be due to the fact that small updates contribute less to training overall, and even when such updates do occur, stochastic rounding may re-adjust them toward the optimal value in the next training steps.
@@ 6 Conclusion
Limitations
While DQT demonstrates strong performance under low-precision and memory-constrained training environments, there are several limitations to be addressed in future work. First, although DQT achieves competitive results compared to high-precision baselines, it has not yet been tested on very large-scale models, e.g., 7B+ models. We plan to extend our approach to such scales when sufficient computational resources become available. In addition, exploring DQT’s applicability to extreme low-precision scenarios (e.g., on hardware that supports true low-bit integer calculations) is also an open direction for future work.
Appendix A Appendix
A.1 Implementation Details
Model configuration.
The detailed configurations for models of different sizes are provided in Table 2. Note that the batch size varies across model sizes during training, but no gradient accumulation is used in any experiment. For each model, the learning rate is selected via grid search over the set {1e-5, 1e-4, 5e-4, 1e-3} using our development set. Regarding the tokenizer, we adopt a publicly released pre-trained one [^5] https://huggingface.co/1bitLLM/bitnet_b1_58-large without further updates during training. We fix the random seed to 42 for reproducibility.
Dataset preprocessing.
The maximum length of training data for both datasets is set to 512. Texts longer than 512 tokens are split into separate chunks, while shorter texts are padded accordingly. For Wikipedia, this preprocessing results in approximately 14 million sentences derived from the original 6.4 million examples, while for FineWeb, the dataset yields around 33 million sentences. The test set of WikiText-2 is from  [^6] https://huggingface.co/datasets/Salesforce/wikitext/viewer/wikitext-2-v1/test.
Low-precision environments.
For FP8 simulation, we choose to use MS-AMP because the GH200 Superchips are equipped with ARM-based CPUs, which are not fully supported by the transformer-engine library currently. As for the reason for using low-precision simulation, hardware and software constraints, such as the inability to modify low-level PyTorch kernels or implement true integer-based weight updates, make it infeasible to perform actual low-bit computation in our environment. These limitations are typical in academic settings, where direct access to low-level hardware accelerators is restricted. As an alternative, we focus on evaluating the practical effectiveness of our method under memory-constrained conditions, simulating the scenarios where computational resources are limited.
Figure 8: An example of our training process using 8-bit quantization.
Params hidden_size intermediate_size num_hidden_layers num_attention_heads batch_size
130M 768 2048 12 12 64
320M 1024 2048 24 16 32
1B 2048 3072 24 32 16
Table 2: Configuration of different models sizes.
A.2 Ternary Inference
Since the straight-through estimator is not employed in our proposed DQT, both the forward and backward operate directly on $n$-bit weight matrices, which means that our DQT models trained with larger bit-width are not inherently ternary during inference. To enable a fair comparison with BitNet b1.58, we adapt our models to perform ternary inference. This is achieved by using ternary weights during the forward pass while maintaining $n$-bit weights for the backward pass, updated through the straight-through estimator. The straight-through estimator is employed only to enable ternary inference in $n$-bit DQT models as a variant of our models. We additionally present the training loss for DQT 8 bit and DQT 8 bit with ternary inference in Figure 9.
Figure 9: Training loss comparison of DQT 8 bits and DQT 8 bits that utilizes ternary inference. DQT achieves ternary inference with minimal degradation.
Figure 10: Non-logarithmic training loss comparison of DQT 8 bits and BitNet b1.58 in 1B sizes. Our DQT 8 bits performs slightly better than BitNet b1.58.
A.3 Actual Memory Usage
We present the actual memory usage of models with different sizes in Table 3. The values reflect usage on a single GH200 GPU, which has 97,871 MB of available memory. For 130M models, we use 4 GPUs, and for 1B models, 16 GPUs are used. We can also observe exactly how much memory is saved after applying Adafactor in BF16 and FP8 for different sizes of models from Table 3.
A.4 Weight Update Frequency
For BitNet models, we compare quantized weight matrices at adjacent training steps by iterating through all corresponding rows and columns to identify whether each element is updated or not. For DQT models, we can directly compare the weight matrices before and after stochastic rounding. The update percentage is computed as the ratio of changed weight elements to the total number of elements.
Model size FP32 BF16 BF16+Adafactor FP8 FP8+Adafactor
130M 69327 54675 53827 39276 38315
1B 76533 58345 53723 40945 37669
Table 3: Actual GPU memory usage (in MB) of models with different sizes on a single GH200.

### v3

@@ 4.5 Evaluation Results
In this section, we conduct evaluation on the WikiText-2 test set (Merity et al., 2016) and several other zero-shot tasks. We focus on 1B parameter models to best demonstrate the capability of our approach. The results are summarized in Table 1. For consistency, we set the sequence length to 512 across all tasks, matching the configuration used during training. In the table, ‘ternary Inf.’ refers to DQT models trained with 8-bit weights but evaluated using ternary weights. Details on how ternary inference is implemented for DQT models are provided in Supplementary Material, Section 3. As shown in Table 1, FP32 models achieve the highest overall performance across tasks, followed closely by our DQT 8-bit variants. Notably, except for the WinoGrande task using 1B models trained on the Wikipedia dataset, our DQT 8 bit models outperform BitNet b1.58 across all other benchmarks. These results highlight that DQT 8 bit models more closely approximate the performance of FP32 models compared to BitNet. Moreover, when ternary inference is applied, the performance slightly decreases compared to 8-bit inference but remains on par with BitNet, demonstrating the robustness of our approach with ternary inference and its flexibility in deployment.
@@ 5 Analysis
The use of stochastic rounding in DQT can be viewed as a convergence-guaranteed optimization scheme. This follows from the fact that stochastic rounding introduces zero-mean noise with bounded variance, a property well studied in stochastic optimization theory (Gupta et al., 2015; Hubara et al., 2018; Bottou et al., 2018; Ajalloeian and Stich, 2021). As a result, the updates in DQT maintain convergence guarantees despite operating in a quantized space. For completeness, we provide a proof sketch of this guarantee in Supplementary Material, Section 4. In this section, we primarily focus on the empirical analysis, demonstrating how these theoretical properties manifest in practice.
While stochastic rounding offers advantages in keeping low-precision weights, it may bring some problems. For example, small weight updates may be applied with low probability (e.g., a value like 1+0.02-1+0.02 has only a 2% chance of being rounded to 0). If such rare updates do occur, we are particularly interested in understanding their impact on the overall training dynamics. In this section, we provide a detailed analysis of the effects and implications of using stochastic rounding.
@@ 5.1 The Role of Stochastic Rounding in DQT
Figure 5: Left: Comparison between DQT 1.58 bits and a variant using absmax quantization for weight updates under the same learning rate. Right: Percentage of updated weights after each training step.
Firstly, we provide empirical evidence for the critical role of stochastic rounding in preserving gradient information during training with low-precision weights. Particularly, we compare 130M ternary DQT with a variant that simply uses absmax quantization on the updated weight matrices and maintains the weights in ternary format without stochastic rounding. As shown in the left part of Figure 5, the latter fails to converge, despite operating under the same bit budget. This outcome is expected, as absmax quantization can easily ignore small updates on the weights. In contrast, stochastic rounding not only performs quantization but also facilitates the accumulation of fine-grained updates, enabling effective training even in extremely low-bit regimes.
@@ 5.2 Quantifying Weight Update Frequency in DQT
We quantify the weight update frequency in the right part of Figure 5 for three variants: DQT 1.58 bit, BitNet b1.58 and DQT 8 bit under the same learning rate and batch size. Note that we show the average percentages of all weight matrices in the model. For BitNet, the peak weight update rate is approximately 0.05% (observed at step 2000, the end of the warm-up phase), meaning only 0.05% of quantized ternary weights change after a single step. Ternary DQT exhibits a similar update rate of around 0.04%, indicating minimal difference between the two in terms of update frequency. In contrast, the 8-bit DQT variant, with weight values ranging from 128-128 to 127, shows a significantly higher update frequency, reaching up to 8%.
@@ 5.3 Impact of Small Weight Updates in DQT Training
Finally, to investigate the impact of small weight updates during training, we rank the absolute value of weight updates at each step and selectively intervene in the lowest 20%. For these bottom 20% updates, we apply one of two interventions: either suppress the update by retaining the original ternary value, or enforce a change by rounding it to a different quantized value, even if the update is small. For example, if an update of +0.02 from 1+0.02-1+0.02 falls within the smallest 20%, we either keep it at 1-1 (suppress) or round it to 0 (enforce), depending on the experimental condition. Figure 6 shows the training loss for these variants, where force to remain refers to suppressing the bottom 20% of updates, and force to update corresponds to enforcing them to change. From Figure 6, we observe that the original implementation of DQT (1.58-bit) achieves the best performance. Ignoring the smallest 20% of updates has minimal impact, while enforcing these small updates to take effect slightly accelerates convergence. However, the final losses at the end of training are similar across all variants. This may be due to the fact that small updates contribute less to training overall, and even when such updates do occur, stochastic rounding may re-adjust them toward the optimal value in the next training steps.